Week of 050912

Open Actions from last week:
  • check that correct SEs are published in prod. BDII. (Patricia)
  • Arrange LFC production meeting with LHCb (Roberto) DONE
  • FTS pilot to check query duration with queue length (Paolo/Gavin) DONE
  • Need to get procedures for dCache (Maarten)
  • LEMON monitors for LFC and LFC-DLI DONE
  • Follow Elonex issue for FTS Pilot DB (Vlado) DONE

  • Need to restart castor2 db to apply patch - schedule for gap between alice and cms (Vlado) SCHEDULED, WAITING 19th SEP
  • Get "dev-kit" for FTS API to write LFC-FPS (Paolo, David) WAIT for FTS1.4
  • FTS Pilot DB Intervention - Thur 15/09 DONE

On Call: Andrea + (Jamie)

Monday:

Log: No Minutes Taken

New Actions:

Discussion:

Tuesday:

Log:

New Actions:

  • RB for CMS (Maarten) DONE
  • FTS1.4 now released - stress test on fts005 with SARA (Gavin) DONE with FZK
  • Problem with LFC-DLI - hit a known bug = check status of alarm deployment (James) DONE

Discussion:

  • Jan VE reported that they were going to move the fts pilot db to another node, but it had hardware problems. Still ongoing - now to do next thursday
  • Kris improved bad query inside FTS that was causing load.
  • Problem with FTS on pre-production system caused machine to run out of processes - fts-support will check.

  • From Maarten
    • CMS now have a second RB, gdrb06
    • YAIM almost ready for the integration of SC3 dCache SEs

Wednesday

Log: castorgridsc needed a config change to work with lcg_utils. Deployed disk tuning fix on wan cluster

Actions:

  • Mail expt's about intervention on pilot next thursday (James). DONE
  • Followup on LFC-DLI lemon problem (James/Jan) DONE

Discussion:

  • Problem with pre-prod was due to a config problem - was running 14 web services, not one. Plan is to move this to another node off devdb anyway - meeting scheduled with Physics DB on Thursday

Thursday

Log:

Actions:

  • Tell ALICE that CMS start on monday (Patricia) DONE
  • Check with FIO about storage for CMS (James) DONE
  • Publish infor in Wiki about FTS servers (Simone) DONE
  • add VO shares to channels on lxshare026d (...)
  • Publish info into WIki about info system (James/Gavin)
  • QF for BDII - and guide for T-2 or expt to set it up (Gavin)

Discussion:

  • ALICE T/F meeting tomorrow. Pablo is on vacation - will start on monday Some jobs submitted.

Friday

Log: problem with deleting via gridftp in castor - ends up with copies on disk - should use SRM - should tell experiments.

Actions:

  • Need to put in caching to BDII provider for FTS (Gavin)

Discussion:

Edit | Attach | Watch | Print version | History: r7 < r6 < r5 < r4 < r3 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r7 - 2006-11-28 - LaurenceField
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback